OpenStudy (anonymous):

I NEED HELP !!Evaluate the integral of x^3/sqrt(25-x^3) dx

OpenStudy (anonymous):

try to substitute x^3 = 25 sin^2y

OpenStudy (anonymous):

Can you please explain why? And are you using trig substitution?

OpenStudy (anonymous):

yes... but not sure if it works or not

OpenStudy (anonymous):

1/3(25-x^3)^(1/2)=x^2 take the first derivative to chk if its the right answer.

OpenStudy (anonymous):

I substitute this ? im so sorry, im really confuse can anyone please show the steps

OpenStudy (anonymous):

thats the evaluated integral but i had to guess the answer and check it with taking the derivative.

zepdrix (zepdrix):

Hmm are you sure the numerator isn't suppose to be x^2? The problem would make a lot more sense... Just wanted to check. Here is ummm... what our solution would look like as the problem stands: http://www.wolframalpha.com/input/?i=integral+x%5E3%2F%5Bsqrt%2825-x%5E3%29%5Ddx Hmm :\